### v1

# Bridging the Sim-to-Real Gap with Bayesian Inference

Abstract
We present Sim-FSVGD for learning robot dynamics from data. As opposed to traditional methods, Sim-FSVGD leverages low-fidelity physical priors, e.g., in the form of simulators, to regularize the training of neural network models. While learning accurate dynamics already in the low data regime, Sim-FSVGD scales and excels also when more data is available. We empirically show that learning with implicit physical priors results in accurate mean model estimation as well as precise uncertainty quantification. We demonstrate the effectiveness of Sim-FSVGD in bridging the sim-to-real gap on a high-performance RC racecar system. Using model-based RL, we demonstrate a highly dynamic parking maneuver with drifting, using less than half the data compared to the state of the art.
I Introduction
For many decades, physical equations of motion have been leveraged to perform highly dynamic and complex tasks in robotics. However, recent advances have shown a significant gap between the physical models and the real world, also known as the ‘sim-to-real’ gap. The gap is primarily because domain-specific models/simulators often tend to neglect complex real-world phenomena (e.g., aerodynamics, system latency, elastic deformation, etc. [1, 2])
Hence, in recent years, data-driven approaches, such as reinforcement learning (RL), have been widely applied in robotics [3]. Particularly, model-based RL methods are often used for robot learning directly on hardware due to their improved sample-efficiency [4, 5, 6, 7]. In model-based deep RL, the dynamics of the real-world system are learned directly from data using neural networks (NNs). However, these methods often ignore the abundance of knowledge available in physical simulators and instead directly fit a model from scratch. We believe that this approach is a critical source of data inefficiency.
In this paper, we discuss how we can harness the benefits of both worlds, i.e., having the flexibility of NNs while retaining the domain knowledge embedded in simulators, encapsulating models of various degrees of fidelity. Moreover, we leverage recent advances in Bayesian inference in functional space [8, 9] to impose a functional prior derived from the simulation/physical model. The prior regularizes the network to behave similarly to the simulation in a low-data regime and fit the real-world dynamics when data is available. Empirically, this results in a substantial sample-efficiency gain.
In summary, our main contributions are; (i) we propose Sim-FSVGD, a simple and tractable update rule for training Baysian Neural Networks (BNNs) that incorporates knowledge for available simulators, (ii) we show that Sim-FSVGD has significant gains in sample-efficiency on learning real-world dynamics with NNs over several other widely applied methods, and (iii) we combine Sim-FSVGD with a model-based RL algorithm and show that it results in considerably faster convergence on a highly dynamic RC Car (cf., fig. 1).
Figure 1: Dynamic RC car from our experiments.
II Related Work.
II-1 Bridging the sim-to-real gap
Methods such as domain randomization, meta-RL, and system identification are typically used to bridge the sim-to-real gap in robotics and RL (cf., [10]). Domain randomization based approaches assume that the real world is captured well within the randomized simulator. In many cases, such as the ones we consider in this work, this is not the case. Meta-RL methods [11, 12, 13, 7] also require that the meta-training and test tasks are i.i.d. samples from a task distribution, and thereby representative. Furthermore, they generally require a plethora of data for meta-training, which is typically also computationally expensive. Instead, Sim-FSVGD does not require any meta training and solely leverages query access to low-fidelity simulators which are based on first-principles physics models and are often available in robotics. Moreover, we leverage solely the real world data and operate in the data-scarce regime. To this end, we consider model-based learning approaches due to their sample efficiency. Typical model-based approaches, perform system identification [14, 15, 16, 17, 18, 19] to fit the simulation parameters on real-world data. However, if the simulation model cannot capture the real-world dynamics well, this results in suboptimal performance (cf.,  section V). Most closely related to our work are approaches such as [2, 20, 21, 22, 23]. They use the simulator model and fit an additional model, e.g., NN or Gaussian process, on top to close the sim-to-real gap. We refer to this approach as GreyBox. In section V we show that our method results in better uncertainty estimates than the GreyBox approach.
II-2 Neural Networks for System Identification
Neural networks are commonly used for learning robot dynamics from data [2, 24, 25, 26, 27, 28] due to their high expressiveness. However, during control, planning with NN models is challenging because most planners exploit (“overfit” to) inaccuracies of the learned model. This leads to suboptimal performance on the real system [29]. BNNs do not suffer from the same pitfall [29]. As a result, recently, BNNs have been widely used for dynamics learning [29, 30, 31, 32, 33, 34].
II-3 Bayesian Neural Networks
Unlike learning a single NN, BNNs maintain a distribution over the NN parameters. Moreover, using a known prior distribution over the NN parameters, given the training data, BNNs infer the posterior distribution. Exact Bayesian inference is computationally intractable for NNs, and approximations [35, 36, 37, 38, 39, 40, 41] are generally used. In this work, we focus on particle-based BNNs due to their computational tractability and reliable uncertainty estimates  [8, 42, 43] (cf., [44] or section III for more details). Nearly all BNNs impose a generic standard normal prior on the NN parameters. On the contrary, we leverage existing (low-fidelity) simulators to impose an informed prior. This boosts sample efficiency when learning robot dynamics from data and also enables bridging the ‘sim-to-real’ gap (cf., section V).
III Background
III-A Learning Dynamics Models with BNNs
The main goal of this work is to efficiently learn a robot dynamics model from data, that is, given a dataset of state-action [^1] Analogous to model-based learning literature, we focus on the fully observable setting. pairs and the next state $({\bm{s}}_{t},{\bm{u}}_{t},{\bm{s}}_{t+1})$, learn a model to predict the next state. To this end, we focus on a general regression problem with a dataset $\mathcal{D}=(\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}})$, comprised of $m$ noisy evaluations $y_{j}=f\left(\mathbf{x}_{j}\right)+\epsilon_{j}$ of an unknown function $f:\mathcal{X}\mapsto\mathcal{Y}$. Here, the training inputs are denoted by $\mathbf{X}^{\mathcal{D}}=\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}$ and corresponding function values by $\mathbf{y}^{\mathcal{D}}=\left\{y_{j}\right\}_{j=1}^{m}$. The noise is assumed to be i.i.d. and Gaussian. To fit a regressor on the data, we employ a NN $h_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ with weights $\theta\in\Theta$. The conditional predictive distribution for the noisy observations can be defined as $p(y|\mathbf{x},\theta)=\mathcal{N}(y|h_{\theta}(\mathbf{x}),\sigma^{2})$ using $h_{\theta}$, with $\sigma^{2}$ as the observation noise variance.
BNNs maintain a distribution over the NN parameters $\theta$. In particular, Bayes’ theorem combines a prior distribution $p(\theta)$ with the empirical data into a posterior distribution $p(\theta|\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}})\propto p(\mathbf{y}^{\mathcal{D}}|\mathbf{X}^{\mathcal{D}},\theta)p(\theta)$. Here, $p(\mathbf{y}^{\mathcal{D}}|\mathbf{X}^{\mathcal{D}},\theta)$ is the likelihood of the data, given a NN hypothesis $\theta$. Under the i.i.d. hypothesis, the likelihood factorizes as $p(\mathbf{y}^{\mathcal{D}}|\mathbf{X}^{\mathcal{D}},\theta)=\prod_{j=1}^{m}p(y_{j}|\mathbf{x}_{j},\theta)$.
To make predictions for an unseen test point $\mathbf{x}^{*}$, we marginalize out the parameters $\theta$ in the posterior, i.e., the predictive distribution is calculated as
[…]
$=\mathbb{E}_{\theta}\left[p(y^{*}|\mathbf{x}^{*},\theta)|\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}}\right].$
III-B BNN Inference in the Function Space
Due to the high-dimensional parameter space $\Theta$ and the over-parameterized nature of the NN mapping $h_{\theta}(\mathbf{x})$, posterior inference for BNNs is challenging. Hence, we need to fall back on some form of approximate inference [45, 46] to make BNN inference tractable. Moreover, the complex parametrization also makes choosing an appropriate prior distribution $p(\theta)$ very hard.
Aiming to alleviate these issues, an alternative approach views BNN inference in the function space, i.e., instead of imposing a prior in the parameter space $\Theta$, a prior in the space of regression functions $h:\mathcal{X}\mapsto\mathcal{Y}$ is used. This results in the posterior $p(h|\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}})\propto p(\mathbf{y}^{\mathcal{D}}|\mathbf{X}^{\mathcal{D}},h)p(h)$ [8, 9]. Here, $p(h)$ is a stochastic process prior, e.g., Gaussian process (GP) [47] with index set $\mathcal{X}$, taking values in $\mathcal{Y}$.
Stochastic processes can be understood as infinite-dimensional random vectors. Hence, they are computationally intractable. However, given finite measurement sets $\mathbf{X}:=[\mathbf{x}_{1},...,\mathbf{x}_{k}]\in\mathcal{X}^{k},k\in\mathbb{N}$, the stochastic process can be characterized by its marginal distributions of function values $\rho(\mathbf{h}^{\mathbf{X}}):=\rho(h(\mathbf{x}_{1}),...h(\mathbf{x}_{k}))$ (cf. Kolmogorov Extension Theorem, [48]). This allows us to break down functional BNN inference into a more tractable form by re-phrasing it in terms of posterior marginals for measurement sets $\mathbf{X}$: $p({\mathbf{h}}^{\mathbf{X}}|\mathbf{X},\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}})\propto p(\mathbf{y}^{\mathcal{D}}|{\mathbf{h}}^{\mathbf{X}^{\mathcal{D}}})p({\mathbf{h}}^{\mathbf{X}})$.
This functional posterior can be tractably approximated for BNN inference [8, 9]. In the following, we briefly describe how this can be done with functional Stein Variational Gradient Descent [FSVGD, 8]: The FSVGD method approximates the posterior as a set of $L$ NN parameter particles $\{\theta_{1},...,\theta_{L}\}$. To improve the particle approximation, FSVGD iteratively re-samples measurement sets $\mathbf{X}$ from a distribution $\nu$ supported on $\mathcal{X}$, e.g., Uniform$(\mathcal{X})$. It then computes SVGD updates [43] in the function space, and projects them back into the parameter space, to update the particle configuration.
To guide the particles towards areas of high posterior probability, FSVGD uses the functional posterior score, i.e., the gradient of the log-density
$\nabla_{{\mathbf{h}}^{\mathbf{X}}}$ $\ln p({\mathbf{h}}^{\mathbf{X}}|\mathbf{X},\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}})=$
$\nabla_{{\mathbf{h}}^{\mathbf{X}}}\ln p(\mathbf{y}^{\mathcal{D}}|{\mathbf{h}}^{\mathbf{X}^{\mathcal{D}}})+\nabla_{{\mathbf{h}}^{\mathbf{X}}}\ln p({\mathbf{h}}^{\mathbf{X}})~{}.$ (1)
In particular, for all $l=1,...,L~{}$ the particle updates are computed as
$\theta^{l}\leftarrow\theta^{l}-\gamma J_{l}u_{l},\quad\text{where}J_{l}=\underbrace{\left(\nabla_{\theta^{l}}{\mathbf{h}}_{\theta^{l}}^{\mathbf{X}}\right)^{\top}}_{\text{NN Jacobian}}$ (2)
$u_{l}=\bigg{(}\underbrace{\frac{1}{L}\sum_{i=1}^{L}\mathbf{K}_{li}\nabla_{{\mathbf{h}}_{\theta^{i}}^{\mathbf{X}}}\ln p({\mathbf{h}}_{\theta^{l}}^{\mathbf{X}}|\mathbf{X},\mathbf{X}^{\mathcal{D}},\mathbf{y}^{\mathcal{D}})+\nabla_{{\mathbf{h}}_{\theta^{l}}^{\mathbf{X}}}\mathbf{K}_{li}}_{\text{SVGD update in the function space}\vspace{-4pt}}\bigg{)},$
where $\mathbf{K}=[k({\mathbf{h}}_{\theta^{l}}^{\mathbf{X}},{\mathbf{h}}_{\theta^{i}}^{\mathbf{X}})]_{li}$ is the kernel matrix between the function values in the measurement points based on a kernel function $k(\cdot,\cdot)$. As we can see in (2), FSVGD only uses the prior scores and, in principle, does not require the density function of the prior marginals. This constitutes a key insight that we will later draw upon in our approach.
IV Sim-FSVGD
In this section, we present Sim-FSVGD. We start this section with an illustrative example to motivate our approach.
Algorithm 1 Sim-FSVGD
Input: Measurement distribution $\nu$, Sim prior
                $g$, Parameter distribution $p(\phi)$, GP
                $p(\tilde{\mathbf{h}})$, Data $\mathcal{D}$, BNN particles $\{\theta_{i}\}^{L}_{i=1}$
1:Sample measurement set $\mathbf{X}\sim\nu$
2:Sample sim finctions $\{\mathbf{h}^{\mathbf{X}}_{i,j}\}^{d_{y}}_{i=0}$ for $j\in\{1,\dots,P\}$.
3:Calculate empirical mean $\{\mathbf{\mu}^{\mathbf{X}}_{i}\}^{d_{y}}_{i=1}$ and variance $\{\mathbf{\Sigma}^{\mathbf{X}}_{i}\}^{d_{y}}_{i=0}$ over the $P$ samples.
4:Approximate Gaussian prior for each $i$, i.e, $p(\mathbf{h}^{\mathbf{X}}_{i})\sim\mathcal{N}(\mathbf{\mu}^{\mathbf{X}}_{i},\mathbf{\Sigma}^{\mathbf{X}}_{i})$
5:Update BNN particles with eq. 2.
IV-A Illustrative Example: the Pendulum
Let us consider a simplified pendulum model where the pendulum’s state $\mathbf{s}=[\theta,\dot{\theta}]$ is comprised of the angle $\theta$ and angular velocity $\dot{\theta}$. We can describe the change of state (i.e., transition function) by the ODE
$\ddot{\theta}=\frac{mgl\sin(\theta)+C_{m}u}{I}~{}.$ (3)
Here $m,l$ and $I$ are the mass, length, and moment of inertia, respectively. The motor at the rotational joint of the pendulum applies the torque $\tau=C_{m}u$ proportional to the control input $u$.
When aiming to accurately predict how the real pendulum system behaves, there are two key sources of uncertainty/error. First, the exact parameters $\phi$ (here, $\phi=[m,l,C_{m},I]$) of the domain-specific model are unknown. However, we can typically narrow down each parameter’s value to a plausible range. This can be captured via a prior distribution $p(\phi)$ over the model parameters. The process of randomly sampling a parameter set $\phi\sim p(\phi)$ and then integrating/solving the ODE in Equation 3 with the corresponding parameters gives random functions. This allows us to implicitly construct a stochastic process of functions that reflect our simplified pendulum model. We call this the domain-model process.
The second source of uncertainty is various physical phenomena such as aerodynamic drag, friction, and motor dynamics that are not captured in (3). This results in a systematic sim-to-real gap. When constructing a prior $p(h)$ for training a BNN model on the real system, we also need to model the sim-to-real gap.
IV-B Main Algorithm
Now we describe our approach, Sim-FSVGD, which combines the domain-model process with a sim-to-real prior to construct an informed functional prior $p(h)$ for learning the dynamics model. The final algorithm is presented in Algorithm 1. First, we summarize how to construct the stochastic process prior.
In practice, we aim to model vector-valued functions, i.e., $h:\mathcal{X}\mapsto\mathcal{Y}$ where $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ and $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$. We factorize the prior over the output dimensions, i.e., $p(h)=\prod_{i=1}^{d_{y}}p(h_{i})$. This allows us to treat each $h_{i}:\mathbf{X}\mapsto\mathbb{R}$ as an independent scalar-valued function.
IV-B1 Domain-model process
The first component of the implicit prior is the domain-specific (low-fidelity) simulation model $g(\mathbf{x},\phi)$ which takes as an input the state action pair $(\mathbf{s},\mathbf{u})$, i.e., $\mathbf{x}=[\mathbf{s},\mathbf{u}]$, as well as system parameters $\phi$. Crucially, we do not require an analytical expression of $g:\mathcal{X}\mapsto\mathcal{Y}$. It is sufficient to have query access, i.e., the output vector of $g(\mathbf{x},\phi)$ can be the result of a numerical simulation (e.g., ODE solver or rigid body simulator).
IV-B2 Sim-to-real prior
The second component is the sim-to-real gap process. For that we employ a GP $p(\tilde{h}_{i})$ per output dimension $i=1,...,d_{y}$ which aims to model the residual/gap between the best possible domain model $g_{i}(\mathbf{x},\phi^{*})$ and the actual target function $f(\mathbf{x})$. We use a GP with zero mean and an isotropic kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\nu^{2}\rho(\left\lVert\mathbf{x}-\mathbf{x}^{\prime}\right\rVert/\ell)$ with variance $\nu^{2}$ and lengthscale $\ell$. Accordingly, the marginal distributions follow a multivariate normal distribution $p(\tilde{\mathbf{h}}^{\mathbf{X}}_{i})=\mathcal{N}(\tilde{\mathbf{h}}^{\mathbf{X}}_{i}|\mathbf{0},\mathbf{K})$ with kernel matrix $\mathbf{K}=[k(\mathbf{x}_{i},\mathbf{x}_{j})]_{i,j}$. The lengthscale $l$ and $\nu^{2}$ are hyperparameters that influence our belief on the sim-to-real gap. A small $\nu^{2}$ implies that the sim-to-real gap is with high probability not too big and a large lengthscale conveys that deviations from the domain model are systematic rather than local.
We combine both processes by independently sampling (conditional) random vectors from both processes and adding them, i.e,
$\begin{split}\mathbf{h}_{i}^{\mathbf{X}}=\hskip 2.0pt&[g_{i}(\mathbf{x}_{1},\phi),...,g_{i}(\mathbf{x}_{k},\phi)]^{\top}+\tilde{\mathbf{h}}^{\mathbf{X}}_{i}\\ &\text{with}~{}~{}\phi\sim p(\phi),~{}\tilde{\mathbf{h}}_{i}^{\mathbf{X}}\sim\mathcal{N}(\tilde{\mathbf{h}}_{i}^{\mathbf{X}}|\mathbf{0},\mathbf{K}).\end{split}$ (4)
The resulting stochastic process prior $p(h)$ is defined implicitly through the marginal distributions that are implied by (4).
IV-B3 Estimating the Stochastic Process Score
To perform FSVGD updates as in (2), we have to compute the score of our stochastic process prior, i.e., $\nabla_{{\mathbf{h}}^{\mathbf{X}}}\ln p({\mathbf{h}}^{\mathbf{X}})=\sum_{i}^{d_{y}}\nabla_{{\mathbf{h}}_{i}^{\mathbf{X}}}\ln p({\mathbf{h}}_{i}^{\mathbf{X}})$. This is intractable for most domain-model processes and we need to estimate it from samples. To this end, we use a simple and efficient Gaussian approximation for the score of prior marginals. In particular, we sample a measurement set $\mathbf{X}$ from a known measurement distribution $\nu$, e.g., uniform distribution over the state-action space. For the measurement set $\mathbf{X}$, we sample $P$ vectors of function values $\mathbf{h}_{i,1}^{\mathbf{X}},...\mathbf{h}_{i,P}^{\mathbf{X}}\sim p(\mathbf{h}_{i}^{\mathbf{X}})$ via (4). Then, we compute their mean and covariance matrix, that is $\mathbf{\mu}^{\mathbf{X}}_{i}$ and $\mathbf{\Sigma}^{\mathbf{X}}_{i}$ and use the score of the corresponding multivariate Gaussian as an approximation, i.e., $p(\mathbf{h}^{\mathbf{X}}_{i})\sim\mathcal{N}(\mathbf{\mu}^{\mathbf{X}}_{i},\mathbf{\Sigma}^{\mathbf{X}}_{i})$.
V Experiments
In this section, we present our experimental results where we evaluate our method, Sim-FSVGD, on learning the robot dynamics from data and online reinforcement learning.
Baselines
We consider three baselines; (i) FSVGD [8], (ii) SysID, and (iii) GreyBox. As discussed in section III, FSVGD is widely applied for Bayesian inference in deep learning. SysID is a simple baseline, that uses the data from the real system to fit the parameters of the simulation model. Crucially, SysID cannot capture the sim-to-real gap. To this end, we also consider GreyBox which first fits the simulation model and then learns a BNN to fit the difference between the real data and the sim model, akin to [20, 22, 23].
V-A One-Dimensional Experiment
Figure 2: BNN posteriors trained on two data points for a one-dimensional sinusoidal function. GreyBox finds accurate mean, but has large uncertainty estimates. At the same time Sim-FSVGD obtains both accurate mean and uncertainty estimates.
We consider a one-dimensional sinusoidal function to visualize the difference between our method and the baselines in fig. 2. Moreover, we train the BNNs with only two data points and plot their posteriors. From fig. 2 we conclude that FSVGD and GreyBox result in very general posteriors that do not capture the sinusoidal nature of the true function. On the contrary, Sim-FSVGD can learn an accurate posterior with only two data points.
Figure 3: We benchmark NLL of dynamics model on test data for Racecar on hardware and in simulations and Pendulum in simulation. BNNs with simulation priors achieve low NLL already in small data regimes and constantly outperform the standard BNNs when we increase the number of train data points.
V-B Sim to Real Transfer
In this experiment, we investigate how Sim-FSVGD bridges the sim-to-real gap for dynamical systems. To this end, we consider a pendulum and racecar in simulation as well as a real-world RC car as our systems. We use an i.i.d. dataset of sampled transitions from the real system to fit a model of the dynamics. As the evaluation metric, we consider the negative log-likelihood (NLL) of our model on a test dataset. For the pendulum, we use the model from eq. 3 as the simulation prior and we simulate the sim-to-real gap by incorporating aerodynamic drag, friction, and motor dynamics in the ‘real’ dynamics. For the racecar, we use the dynamics model from [49] which models tire dynamics with the Pacejka tire model [50]. For the simulation prior, we use the kinematics bicycle model also from [49], which does not capture the tire dynamics. We use the same simulation prior for real RC car. The car consists of a high-torque motor, which allows us to perform dynamic maneuvers that involve loss of traction and drifting. These behaviors are neither captured by our simulation prior nor the dynamics model from [49].
[…]
Our results are presented in fig. 3. From the figure, we conclude that Sim-FSVGD consistently outperforms all other methods in both case studies. Moreover, in the low-data regime, our method achieves orders of magnitude lower NLL, even though there is a considerable sim-to-real gap (c.f., SysID). Furthermore, when more data is provided our method performs similarly to other baselines. Therefore, Sim-FSVGD relies less on the simulation prior when more data is available. This empirically validates the consistency of our approach. In fig. 3, we also study the effect of incorporating the additional sim-to-real GP prior in Sim-FSVGD. We conclude that including the additional GP leads to better performance and thus plays a crucial role, particularly for the pendulum and racecar.
V-C Online RL
[…]
Refer to captionpt
Sim-FSVGD
pt
FSVGD
[…]
Figure 5: Realized trajectory for Racecar environment in simulation after episode 3. In violet, we have the trajectory of the model with FSVGD (no prior), and in red the trajectory of Sim-FSVGD using the low-fidelity prior. We observe that, Sim-FSVGD already learns to get close to the target in three episodes, whereas FSVGD (no prior) explores away from it.
Refer to captionpt
Sim-FSVGD
pt
FSVGD
ptEpisode 1ptTargetptEpisode 1ptTargetptEpisode 10ptTargetptEpisode 10ptTarget
Figure 6: Realized trajectories after episodes one and ten for Sim-FSVGD and FSVGD. Sim-FSVGD moves close to the target in episode one and parks nearly perfectly in episode ten. On the contrary, FSVGD fails to move at all in the first episode and explores away from the target in the tenth.
Figure 7: The simulation prior helps initially by exploring more directly towards the target. Both in simulation and on hardware the agent solves the task significantly faster and with less collected data.
We study the problem of model-based reinforcement learning for the racecar in simulation and the real RC car. In model-based RL, BNNs with good uncertainty estimates play a crucial role in exploration [29, 30, 31, 32, 33]. Moreover, incorporating a simulation prior can lead to directed exploration and therefore faster convergence, i.e., better sample efficiency. To this end, in this experiment, we compare the performance of Sim-FSVGD and FSVGD in model-based RL. Moreover, we use our BNN to obtain a policy $\pi$, which we rollout on the real system to get further data for updating our model (see algorithm 2). We repeat this procedure for 20 episodes and consider a rollout length of $100$. For the policy $\pi$, we use the SAC algorithm [51] which we train using our BNN similar to [52].
Algorithm 2 Sim-FSVGD for Model-Based RL
Input: Initial model and policy ${\bm{f}}_{\Theta}$, $\pi_{\psi}$, $\mathcal{D}:=\emptyset$
1:for $\textnormal{episode}=1,2,\ldots$ do
2:     ${\bm{f}}_{\Theta}\leftarrow\text{{Sim-FSVGD}}({\bm{f}}_{\Theta},\mathcal{D})$
3:     $\pi_{\psi}\leftarrow\text{SAC}(\pi_{\psi},{\bm{f}}_{\Theta})$
4:     $(s_{0},a_{0},...,a_{T-1},s_{T})\leftarrow\textsc{Rollout}(\pi_{\psi})$
5:     $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t},a_{t},s_{t+1})\}_{t=0}^{T-1}$
6:end for
[…]
In Figure 7 we report the learning curves for the racecar in simulation and the real RC car. We observe that in both cases, incorporating the simulation prior leads to considerably faster convergence (ca. 2×2$\times$2 × fewer episodes). Moreover, in Figure 3 we show the realized trajectory for the third episode of the racecar. We see that already after three episodes, Sim-FSVGD learns to drive the car close to the target whereas not incorporating the prior leads to exploration away from the target position. Similar behavior is observed on hardware (c.f., Figure 6), where from episode 1, the policy learned using Sim-FSVGD drives the car close to the target and at episode 10 it is parking the car nearly perfectly. On the contrary, not incorporating the prior leads to no movement of the car at episode 1 and the car driving away from the target at episode 10. This experiment demonstrates that incorporating a simulation prior in model-based RL leads to directed exploration which results in better sample-efficiency.
VI Conclusion
In this work, we address the sim-to-real gap in robotics by harnessing both domain knowledge and data. Leveraging ideas from Bayesian inference in functional space, we propose Sim-FSVGD, a simple and tractable algorithm to incorporate physical simulation priors in learning neural network robot dynamics. Our resulting algorithm benefits from both the expressiveness of neural networks and the domain knowledge from physical simulators. We show that Sim-FSVGD results in significantly better sample efficiency for supervised learning of robot dynamics and model-based RL.
Future Work. Here, we use a simple Gaussian approximation for score estimation. There are a range of other score estimation techniques [e.g. 53, 54]. We leave a detailed comparison among the different methods as future work. Moreover, we regard applying Sim-FSVGD to more high-dimensional systems from robotics or other scientific domains as an interesting avenue for future work.
VII Acknowledgments
This project has received funding from the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and the Microsoft Swiss Joint Research Center. Jonas Rothfuss was supported by an Apple Scholars in AI/ML fellowship.